# No More Hidden Pitfalls? Exposing Smart Contract Bad Practices with LLM-Powered Hybrid Analysis

### v1

@@ 1. Introduction
With the widespread use of blockchain technology, smart contracts have become an important part of the blockchain ecosystem (Sharma_2023_review; huang2025comparative). Smart contracts are computer programs that automatically execute contract terms, controlling assets and operations on the chain. However, due to their public and immutable code, smart contracts have become a significant target for attackers (barboni2022smart). A total of 464 security incidents occurred in 2023, resulting in losses of up to \$2.486 billion (slowmist). The most significant attack occurred on September 23rd when Mixin Network’s cloud service provider database was attacked, involving approximately \$200 million.
@@
Bad practices refer to poor coding habits or design decisions in the development of smart contracts (reyes2023continuous; ding2025comprehensive; wu2025security). We categorize bad practices into two types: (1) Security-related bad practices, which include actual vulnerabilities that may lead to security incidents (e.g., reentrancy); and (2) Quality-related bad practices, which encompass code quality issues that affect maintainability, efficiency, and design quality without directly resulting in exploitable vulnerabilities (e.g., code duplication). While security-related bad practices could potentially lead to future security threats, both types have significant impacts: they can cause performance problems, increase security risks, lead to unpredictable code behavior (10.1145/3711901), and create hidden economic dangers due to the disruption of regular smart contract activities (sharma2023mixed).
@@
Currently, the security audit of smart contracts mainly relies on manual code review and automated tools (10.1145/3641846; yang2025multi). However, these methods have their limitations (10.1145/3674725; 10.1145/3628160; li2025penetrating). Manual code review is inefficient and prone to overlook subtle security vulnerabilities. Existing automated tools primarily rely on pattern matching, which cannot accurately detect complex security issues. Moreover, the types of vulnerabilities that these tools can detect are usually relatively limited and may not be able to identify all potential security problems in smart contracts (chaliasos2024smart). To achieve a comprehensive audit, multiple tools may be required, each covering different aspects of security. Therefore, effectively detecting and preventing security issues in smart contracts remains an important issue that needs to be solved.
@@
To address these challenges, we propose SCALM (Smart Contract Audit Language Model), a framework with two innovations for smart contract auditing. First, a static analysis module constructs an extensible knowledge base by extracting code patterns, converting them into semantic vectors via an embedding model, and dynamically updating the repository with newly identified bad practices through automated SWC-ID annotation. Second, a multi-layer reasoning verification system combines Retrieval-Augmented Generation (RAG) with Step-Back prompting, enabling hierarchical reasoning from syntax checks to architectural risk analysis. This framework detects both explicit vulnerabilities and latent design flaws, generating structured audit reports. This approach generates detailed audit reports that document identified issues, assess risk scores, and provide concrete remediation suggestions, achieving improved detection accuracy compared to existing tools while maintaining practical applicability for developers. Our contributions are as follows:
@@ 2.1. Large Language Models
LLMs are trained using deep learning techniques to understand and generate human language. They are typically based on the Transformer architecture, such as ChatGPT (Kasneci_2023_ChatGPT), BERT (Devlin_2019_BERT), and GLM (Du_2022_GLM). The training process usually involves learning language patterns and structures from large-scale text corpora. These corpora can include a variety of texts such as news articles, books, web pages, and other forms of human linguistic expression. These model can generate coherent and meaningful text by learning from these corpora. Furthermore, LLMs can handle various natural language processing tasks, including text generation, text classification, sentiment analysis, question-answering systems, etc (10.1145/3709358; li2025ckg; li2025facial).
One of the key features of LLMs is their powerful generative ability. These models can generate new, coherent text similar in grammar and semantics to the training data (wu2023autogen). This makes LLMs useful for various applications, including machine translation, text summarization, sentiment analysis, dialogue systems, and other natural language processing tasks (10.1145/3699597). Another important feature of LLMs is their ”zero-shot” capability, which allows them to perform various tasks without any task-specific training (abbasiantaeb2024let). For example, the model can choose the most appropriate answer given a question and some answer options. This ability makes LLMs very useful in many practical applications.
However, LLMs also have some challenges and limitations. For instance, they may generate inaccurate or misleading information and reflect biases in the training data (YAO2024100211). LLM can adopt the RAG to improve quality and accuracy. This method combines pre-trained parametric models with non-parametric memory to enhance the quality and accuracy of smart contract code audits (Gao_2024_Retrieval-Augmented). The RAG integrates the processes of retrieval and generation into one.
As Fig. 1 illustrates, during the operation of the model, it first retrieves relevant documents or entities from a large-scale knowledge base. Then, it inputs this retrieved information as additional context into the generation model, which generates corresponding outputs based on these inputs. This design allows RAG to utilize external knowledge bases effectively while demonstrating excellent performance when dealing with tasks requiring extensive background knowledge (dos2024domain).
@@ 2.2. Smart Contract Weakness Classification
Smart contracts are self-executing protocols that run on the blockchain and allow trusted transactions without third-party intervention (gao2025implementation; liu2025empirical). However, since their code is publicly available and cannot be changed once deployed, the security of smart contracts has become an important issue (10123449). To address this issue, EIP-1470 (wagner2018eip) proposes the Smart Contract Weakness Classification (SWC), a classification scheme designed to help developers identify and prevent smart contract weaknesses.
SWC concerns weaknesses that can be identified within a smart contract’s Solidity code (soud2024fly; li2025uscsa; li2025atomgraph). It is designed to reference the structure and terminology of the Common Weakness Enumeration (CWE) but adds several weakness classifications specific to smart contracts (chen2024healthier). These classifications include but are not limited to, reentry attacks, arithmetic overflow, and delegatecall to untrusted callee. The example in Fig. 2 illustrates how SWC guides secure design: the vulnerable Proxy contract allows any caller to invoke delegatecall with arbitrary addresses, while the fixed version restricts the callee address to an owner-controlled parameter, mitigating unauthorized code execution risks.
@@ 3. Method
Fig. 3 shows the architecture of SCALM, which establishes a systematic framework for detecting bad practices in smart contracts through multi-layer reasoning and verification. SCALM consists of two core modules : (1) To statically analyze and extract bad practice patterns, convert them into semantic vectors, and construct an extensible knowledge base. (2) Combines RAG and Step-Back prompting in a multi-layer reasoning verification system, which realizes multi-stage verification through layered abstraction from code syntax to architecture. Ultimately, SCALM leverages LLMs to generate structured audit reports.
@@ 3.2. Semantic Vectorization and Retrieval
We use a vector database to store and query large amounts of vector data. A vector database is a particular database that can store and query large amounts of vector data (Hambardzumyan_2022_Deepa). In the vector database, data is stored as vectors, each typically represented by a set of floating-point numbers. These vectors can represent various data types, such as images, audio, text, etc. In bad practice detection tasks, through the embedding model (i.e., text-embedding-ada-002), code slices are transformed into 1536-dimensional vectors that preserve syntactic structures and semantic relationships. This process captures both explicit bad practice patterns and implicit code quality issues. This process can be expressed with the following eq. 3:
@@ 3.3. Multi-layer Reasoning and Verification
We employ Step-Back prompting (Zheng__TAKE) to bridge concrete code patterns with abstract security principles. This technique uses the capabilities of LLMs to abstract high-level concepts and basic principles from specific code instances. In this way, not only can the model understand the literal meaning of the code, but it can also comprehend underlying logic and potential design patterns through abstract thinking. Step-Back prompting consists of two main steps:
@@ 3.3.1. Compliance Validation Pipeline
The formula given by eq. 6 calculates a comprehensive risk score by averaging findings across three distinct validation stages ($i=1$ to $3$), each corresponding to a specific security assessment layer. For each layer, the detected bad practices are assigned a normalized severity value derived from CVSS v3.1 baselines (cvss).
@@ 4.1. Experiment Settings
Dataset. In this paper, we use the DAppSCAN dataset (zheng2023dappscan) as a knowledge base for detecting bad practices in smart contracts. The dataset contains 39,904 Solidity files with 1,618 SWC weaknesses from 682 real projects. The Smartbugs dataset (DurieuxEtAl2020ICSE) is also used in the experiment, and a total of 1,894 smart contracts with five types of security-related bad practices (SWC weaknesses) are extracted for comparison experiments. Additionally, we introduce the SolQuality dataset (solquality), comprising 1,200 smart contracts covering six categories of quality-related bad practices. This dataset encompasses 25 diverse contract scenarios, including ERC20 tokens, NFTs (li2025beyond), voting systems, auctions, and other common decentralized application patterns (long2025fomo3d; sun2025data), providing comprehensive coverage for quality-oriented bad practices detection. These datasets form the basis of our experimental analysis. Table 1 summarizes the smart contract data used.
Models. For the selection of LLMs, we chose six current state-of-the-art models for detection experiments. GPT-4o, GPT-4-1106-preview, and GPT-4-0409 are the latest versions from OpenAI with powerful natural language processing capabilities (achiam2023gpt). Claude-3.5-Sonnet is Anthropic’s new-generation model focused on safety and interoperability (jiang2025deepseek). Gemini-1.5-Pro is Google’s high-performance model optimized for multitasking (team2024gemini). Llama-3.1-70b-Instruct is Meta’s large-scale model specializing in instruction following and generating high-quality text (grattafiori2024llama).
@@ 4.4. RQ3: Comparison Experiments
Additionally, we collect a set of smart contract defect detection tools from reputable journals and conferences in software and security (e.g., CCS and ASE) as well as Mythril (mythril), recommended by the official Ethereum community. For comparative analysis, we choose seven benchmark smart contract detection tools: four traditional tools (Mythril, Oyente (Luu_2016_Making), Confuzzius (Torres_2021_ConFuzzius), and Conkas (VelosoConkas)) and three LLM-based tools (GPTLens (hu2023large), VulnHunt-GPT (Boi_2024_VulnHunt-GPT), and LLM-SmartAudit (wei2025advanced)). Several factors are considered in the selection of the tools: (1) The accessibility of the tool’s source code. (2) The tool’s ability to detect the five categories of bad practices we select. (3) The tool’s support for source code written in Solidity. (4) The tool’s ability to report the exact location of potentially defective code for manual review. Note that traditional tools (Mythril, Oyente, Confuzzius, Conkas) focus exclusively on security-related bad practices and cannot evaluate quality-related aspects, while LLM-based tools support both dimensions.
@@ 5. Discussion
Despite its strengths, SCALM has limitations in understanding blockchain mechanisms and complex interactions (zhang2025attacks; zhang2025security). Specifically, SCALM might not fully grasp the state-dependent nature of certain vulnerabilities, which require a deep contextual understanding of how the contract evolves over time. Future work will focus on improving SCALM by incorporating advanced LLMs and fine-tuning with domain-specific data. Additionally, we plan to enhance the RAG generation strategy to provide more prosperous and accurate contextual information. These improvements will improve the performance of SCALM for smart contract auditing.
@@ 6. Related Work
LLMs have been widely applied and validated for their ability to identify and fix vulnerabilities (Napoli_2023_Evaluating; 10.1145/3643674). In the field of smart contract security, various methods based on LLMs have been proposed, and certain effects have been achieved. Firstly, Boi et al. (Boi_2024_VulnHunt-GPT) proposed VulnHunt-GPT, a method that uses GPT-3 to identify common vulnerabilities in OWASP-standard smart contracts. Building on this, Boi et al. (10.1145/3687251.3687253) further demonstrated how LLMs trained on diverse vulnerability datasets can detect multiple security flaws simultaneously, outperforming traditional static analysis tools. Similarly, Xia et al. (Xia_2024_AuditGPT) introduced AuditGPT, which utilizes LLMs to verify ERC rules through decomposed audit tasks. In terms of fuzz testing, Shou et al. (Shou_2024_LLM4Fuzz) proposed LLM4Fuzz to guide fuzz testing priorities, while Zhao et al. (zhao2025detecting) enhanced functional bug detection by combining LLM-driven vulnerability pinpointing with bug-oriented fuzzing. Ding et al. (DING2025126479) developed SmartGuard, leveraging LLMs for semantic code retrieval and Chain-of-Thought reasoning to achieve 95% F1-scores on benchmarks. Wei et al. (wei2024llm) advanced multi-agent collaborative auditing with LLM-SmartAudit, detecting complex logic vulnerabilities overlooked by traditional tools.
However, the direct use of pre-trained LLMs is no longer sufficient on some occasions, so many studies have chosen to fine-tune LLMs to meet specific needs. For example, Luo et al. (10664408) proposed FELLMVP, an ensemble framework that fine-tunes eight LLMs for specific vulnerabilities and integrates their predictions, achieving 98.8% accuracy. Mothukuri et al. (10664261) introduced LLMSmartSec, which fine-tunes GPT-4 to analyze contracts from developer, auditor, and ethical hacker perspectives, then distills knowledge into an LLMGraphAgent via annotated control flow graphs. Liu et al. (Liu_2024_PropertyGPT) designed PropertyGPT for formal verification by generating contract properties through retrieval-augmented LLMs. Storhaug et al. (Storhaug_2023_Efficient) reduced vulnerable code generation via vulnerability-bound decoding, while Yang et al. (Yang_2024_Automated) fine-tuned Llama-2-13B for DeFi-specific vulnerability detection. In conclusion, while LLMs show great potential for improving smart contract security, the full realization of their potential requires continuous fine-tuning and adaptation of these models in specific contexts.

### v2

@@ 1. Introduction
With the widespread use of blockchain technology, smart contracts have become an important part of the blockchain ecosystem (Sharma et al., 2023a; Huang et al., 2025; Cai et al., 2024). Smart contracts are computer programs that automatically execute contract terms, controlling assets and operations on the chain (Han et al., 2026). However, due to their public and immutable code, smart contracts have become a significant target for attackers (Barboni et al., 2022). A total of 464 security incidents occurred in 2023, resulting in losses of up to \$2.486 billion (Zone, 2024). The most significant attack occurred on September 23rd when Mixin Network’s cloud service provider database was attacked, involving approximately \$200 million.
@@
Bad practices refer to poor coding habits or design decisions in the development of smart contracts (Reyes et al., 2023; Ding et al., 2025b; Wu et al., 2025). We categorize bad practices into two types: (1) Security-related bad practices, which include actual vulnerabilities that may lead to security incidents (e.g., reentrancy); and (2) Quality-related bad practices, which encompass code quality issues that affect maintainability, efficiency, and design quality without directly resulting in exploitable vulnerabilities (e.g., code duplication). While security-related bad practices could potentially lead to future security threats, both types have significant impacts: they can cause performance problems, increase security risks, lead to unpredictable code behavior (Kim et al., 2025), and create hidden economic dangers due to the disruption of regular smart contract activities (Sharma et al., 2023b).
@@
Currently, the security audit of smart contracts mainly relies on manual code review and automated tools (Gao et al., 2024a; Yang et al., 2025; Wang and Zhao, 2024; Grossman et al., 2024). However, these methods have their limitations (Ye et al., 2024; Zhou et al., 2023b; Li et al., 2025c; Ayub et al., 2024). Manual code review is inefficient and prone to overlook subtle security vulnerabilities. Existing automated tools primarily rely on pattern matching, which cannot accurately detect complex security issues. Moreover, the types of vulnerabilities that these tools can detect are usually relatively limited and may not be able to identify all potential security problems in smart contracts (Chaliasos et al., 2024). To achieve a comprehensive audit, multiple tools may be required, each covering different aspects of security. Therefore, effectively detecting and preventing security issues in smart contracts remains an important issue that needs to be solved.
@@
To address these challenges, we propose SCALM (Smart Contract Audit Language Model), a framework with two innovations for smart contract auditing (Chen et al., 2024b). First, a static analysis module constructs an extensible knowledge base by extracting code patterns, converting them into semantic vectors via an embedding model, and dynamically updating the repository with newly identified bad practices through automated SWC-ID annotation (Aguilar et al., 2024; Arceri et al., 2024). Second, a multi-layer reasoning verification system combines Retrieval-Augmented Generation (RAG) with Step-Back prompting, enabling hierarchical reasoning from syntax checks to architectural risk analysis. This framework detects both explicit vulnerabilities and latent design flaws, generating structured audit reports. This approach generates detailed audit reports that document identified issues, assess risk scores, and provide concrete remediation suggestions, achieving improved detection accuracy compared to existing tools while maintaining practical applicability for developers. Our contributions are as follows:
@@ 2.1. Large Language Models
LLMs are trained using deep learning techniques to understand and generate human language. They are typically based on the Transformer architecture, such as ChatGPT (Kasneci et al., 2023), BERT (Devlin et al., 2019), and GLM (Du et al., 2022). The training process usually involves learning language patterns and structures from large-scale text corpora. These corpora can include a variety of texts such as news articles, books, web pages, and other forms of human linguistic expression. These model can generate coherent and meaningful text by learning from these corpora. Furthermore, LLMs can handle various natural language processing tasks, including text generation, text classification, sentiment analysis, question-answering systems, etc (Fan et al., 2025; Li et al., 2025b, f).
One of the key features of LLMs is their powerful generative ability. These models can generate new, coherent text similar in grammar and semantics to the training data (Wu et al., 2023). This makes LLMs useful for various applications, including machine translation, text summarization, sentiment analysis, dialogue systems, and other natural language processing tasks (Xiang et al., 2025). Another important feature of LLMs is their ”zero-shot” capability, which allows them to perform various tasks without any task-specific training (Abbasiantaeb et al., 2024). For example, the model can choose the most appropriate answer given a question and some answer options. This ability makes LLMs very useful in many practical applications.
However, LLMs also have some challenges and limitations. For instance, they may generate inaccurate or misleading information and reflect biases in the training data (Yao et al., 2024). LLM can adopt the RAG to improve quality and accuracy. This method combines pre-trained parametric models with non-parametric memory to enhance the quality and accuracy of smart contract code audits (Gao et al., 2024b). The RAG integrates the processes of retrieval and generation into one.
As Fig. 1 illustrates, during the operation of the model, it first retrieves relevant documents or entities from a large-scale knowledge base. Then, it inputs this retrieved information as additional context into the generation model, which generates corresponding outputs based on these inputs. This design allows RAG to utilize external knowledge bases effectively while demonstrating excellent performance when dealing with tasks requiring extensive background knowledge (dos Santos Junior et al., 2024).
@@ 2.2. Smart Contract Weakness Classification
Smart contracts are self-executing protocols that run on the blockchain and allow trusted transactions without third-party intervention (Gao et al., 2025; Liu et al., 2025). However, since their code is publicly available and cannot be changed once deployed, the security of smart contracts has become an important issue (Zhou et al., 2023a). To address this issue, EIP-1470 (Wagner, 2018) proposes the Smart Contract Weakness Classification (SWC), a classification scheme designed to help developers identify and prevent smart contract weaknesses.
SWC concerns weaknesses that can be identified within a smart contract’s Solidity code (Soud et al., 2024; Li et al., 2025e, d). It is designed to reference the structure and terminology of the Common Weakness Enumeration (CWE) but adds several weakness classifications specific to smart contracts (Chen et al., 2024a). These classifications include but are not limited to, reentry attacks, arithmetic overflow, and delegatecall to untrusted callee. The example in Fig. 2 illustrates how SWC guides secure design: the vulnerable Proxy contract allows any caller to invoke delegatecall with arbitrary addresses, while the fixed version restricts the callee address to an owner-controlled parameter, mitigating unauthorized code execution risks.
@@ 3. Method
Fig. 3 shows the architecture of SCALM, which establishes a systematic framework for detecting bad practices in smart contracts through multi-layer reasoning and verification (Sun et al., 2025b; He et al., 2024). SCALM consists of two core modules : (1) To statically analyze and extract bad practice patterns, convert them into semantic vectors, and construct an extensible knowledge base. (2) Combines RAG and Step-Back prompting in a multi-layer reasoning verification system, which realizes multi-stage verification through layered abstraction from code syntax to architecture. Ultimately, SCALM leverages LLMs to generate structured audit reports.
@@ 3.2. Semantic Vectorization and Retrieval
We use a vector database to store and query large amounts of vector data. A vector database is a particular database that can store and query large amounts of vector data (Hambardzumyan et al., 2022). In the vector database, data is stored as vectors, each typically represented by a set of floating-point numbers. These vectors can represent various data types, such as images, audio, text, etc. In bad practice detection tasks, through the embedding model (i.e., text-embedding-ada-002), code slices are transformed into 1536-dimensional vectors that preserve syntactic structures and semantic relationships. This process captures both explicit bad practice patterns and implicit code quality issues. This process can be expressed with the following eq. 3:
@@ 3.3. Multi-layer Reasoning and Verification
We employ Step-Back prompting (Zheng et al., 2024a) to bridge concrete code patterns with abstract security principles. This technique uses the capabilities of LLMs to abstract high-level concepts and basic principles from specific code instances. In this way, not only can the model understand the literal meaning of the code, but it can also comprehend underlying logic and potential design patterns through abstract thinking. Step-Back prompting consists of two main steps:
@@ 3.3.1. Compliance Validation Pipeline
The formula given by eq. 6 calculates a comprehensive risk score by averaging findings across three distinct validation stages ($i=1$ to $3$), each corresponding to a specific security assessment layer. For each layer, the detected bad practices are assigned a normalized severity value derived from CVSS v3.1 baselines (FIRST, 2019).
@@ 4.1. Experiment Settings
Dataset. In this paper, we use the DAppSCAN dataset (Zheng et al., 2024b) as a knowledge base for detecting bad practices in smart contracts. The dataset contains 39,904 Solidity files with 1,618 SWC weaknesses from 682 real projects. The Smartbugs dataset (Durieux et al., 2020) is also used in the experiment, and a total of 1,894 smart contracts with five types of security-related bad practices (SWC weaknesses) are extracted for comparison experiments. Additionally, we introduce the SolQuality dataset (RikkaLzw, 2025), comprising 1,200 smart contracts covering six categories of quality-related bad practices. This dataset encompasses 25 diverse contract scenarios, including ERC20 tokens, NFTs (Li et al., 2025a), voting systems, auctions, and other common decentralized application patterns (Long et al., 2025; Sun et al., 2025a), providing comprehensive coverage for quality-oriented bad practices detection. These datasets form the basis of our experimental analysis. Table 1 summarizes the smart contract data used.
Models. For the selection of LLMs, we chose six current state-of-the-art models for detection experiments. GPT-4o, GPT-4-1106-preview, and GPT-4-0409 are the latest versions from OpenAI with powerful natural language processing capabilities (Achiam et al., 2023). Claude-3.5-Sonnet is Anthropic’s new-generation model focused on safety and interoperability (Jiang et al., 2025). Gemini-1.5-Pro is Google’s high-performance model optimized for multitasking (Team et al., 2024). Llama-3.1-70b-Instruct is Meta’s large-scale model specializing in instruction following and generating high-quality text (Grattafiori et al., 2024).
@@ 4.4. RQ3: Comparison Experiments
Additionally, we collect a set of smart contract defect detection tools from reputable journals and conferences in software and security (e.g., CCS and ASE) as well as Mythril (Mythril, 2024), recommended by the official Ethereum community. For comparative analysis, we choose seven benchmark smart contract detection tools: four traditional tools (Mythril, Oyente (Luu et al., 2016), Confuzzius (Torres et al., 2021), and Conkas (Veloso, 2024)) and three LLM-based tools (GPTLens (Hu et al., 2023), VulnHunt-GPT (Boi et al., 2024b), and LLM-SmartAudit (Wei et al., 2025)). Several factors are considered in the selection of the tools: (1) The accessibility of the tool’s source code. (2) The tool’s ability to detect the five categories of bad practices we select. (3) The tool’s support for source code written in Solidity. (4) The tool’s ability to report the exact location of potentially defective code for manual review. Note that traditional tools (Mythril, Oyente, Confuzzius, Conkas) focus exclusively on security-related bad practices and cannot evaluate quality-related aspects, while LLM-based tools support both dimensions.
@@ 5. Discussion
Despite its strengths, SCALM has limitations in understanding blockchain mechanisms and complex interactions (Zhang et al., 2025a, b). Specifically, SCALM might not fully grasp the state-dependent nature of certain vulnerabilities, which require a deep contextual understanding of how the contract evolves over time. Future work will focus on improving SCALM by incorporating advanced LLMs and fine-tuning with domain-specific data. Additionally, we plan to enhance the RAG generation strategy to provide more prosperous and accurate contextual information. These improvements will improve the performance of SCALM for smart contract auditing.
@@ 6. Related Work
LLMs have been widely applied and validated for their ability to identify and fix vulnerabilities (Napoli and Gatteschi, 2023; Liu et al., 2024b). In the field of smart contract security, various methods based on LLMs have been proposed, and certain effects have been achieved. Firstly, Boi et al. (Boi et al., 2024b) proposed VulnHunt-GPT, a method that uses GPT-3 to identify common vulnerabilities in OWASP-standard smart contracts. Building on this, Boi et al. (Boi et al., 2024a) further demonstrated how LLMs trained on diverse vulnerability datasets can detect multiple security flaws simultaneously, outperforming traditional static analysis tools. Similarly, Xia et al. (Xia et al., 2024) introduced AuditGPT, which utilizes LLMs to verify ERC rules through decomposed audit tasks. In terms of fuzz testing, Shou et al. (Shou et al., 2024) proposed LLM4Fuzz to guide fuzz testing priorities, while Zhao et al. (Zhao et al., 2025) enhanced functional bug detection by combining LLM-driven vulnerability pinpointing with bug-oriented fuzzing. Ding et al. (Ding et al., 2025a) developed SmartGuard, leveraging LLMs for semantic code retrieval and Chain-of-Thought reasoning to achieve 95% F1-scores on benchmarks. Wei et al. (Wei et al., 2024) advanced multi-agent collaborative auditing with LLM-SmartAudit, detecting complex logic vulnerabilities overlooked by traditional tools.
However, the direct use of pre-trained LLMs is no longer sufficient on some occasions, so many studies have chosen to fine-tune LLMs to meet specific needs. For example, Luo et al. (Luo et al., 2024) proposed FELLMVP, an ensemble framework that fine-tunes eight LLMs for specific vulnerabilities and integrates their predictions, achieving 98.8% accuracy. Mothukuri et al. (Mothukuri et al., 2024) introduced LLMSmartSec, which fine-tunes GPT-4 to analyze contracts from developer, auditor, and ethical hacker perspectives, then distills knowledge into an LLMGraphAgent via annotated control flow graphs. Liu et al. (Liu et al., 2024a) designed PropertyGPT for formal verification by generating contract properties through retrieval-augmented LLMs. Storhaug et al. (Storhaug et al., 2023) reduced vulnerable code generation via vulnerability-bound decoding, while Yang et al. (Yang et al., 2024) fine-tuned Llama-2-13B for DeFi-specific vulnerability detection. In conclusion, while LLMs show great potential for improving smart contract security, the full realization of their potential requires continuous fine-tuning and adaptation of these models in specific contexts.